# Elaborating Higgs to dimuon decay from gluon fusion by decorrelation and jet substructure

### v1

@@ III.1 Comparison
As mentioned before, the decorrelation is in conflict with discrimination. Therefore the comparison among three machine discrimminators - BDT, (ordinary)DNN and decorrelated DNN - can be summarized in terms of two properties, the performance of discrimination and the degree of decorrelation.
Figure 4: Receiver operating characteristic(ROC) curve for three discrimination schemes - BDT, (ordinary) DNN and decorrelated DNN - with two different set of variables($S_{n_{\text{jet}}}$ and $JS_{n_{\text{jet}}})$. The area under curve(AUC) is also given. Note that the true positive rate is equivalent to signal efficiency and the false positive rate is equivalent to (1-background rejection rate).
Fig. 4 shows the receiver operating characteristic (ROC) curve with area under curve (AUC). It shows the performance of each discriminator. For all three discriminators, including jet substructure variables (solid lines) improve the discrimination performance of ggH channel. The ordinary DNN’s (black) have large AUC since they exploited $m_{\mu\mu}$ internally. BDT (blue) and decorrelated DNN (red) have similar AUCs, which is around 0.1 less than those of ordinary DNN. The gap between ordinary DNN’s and other methods represents the discrimination power of $m_{\mu\mu}$, or more conservatively, power of the information of training variables correlated to $m_{\mu\mu}$. The small AUC’s of decorrelated DNN reflects that the decorrelation intentionally discards the information about $m_{\mu\mu}$ and those of BDT reflects that the BDT cannot reconstruct $m_{\mu\mu}$ internally. The small difference of AUC between BDT and decorrelated can be interpreted as the the uncontrolled contamination by $m_{\mu\mu}$.
The degree of decorrelation between $m_{\mu\mu}$ and the output of discriminators can be seen directly from the distortion of invariant mass distribution categorized by the output of discriminators (Fig.1, 2, 3). It can also be quantitatively checked by a test of independence using asymptotic distribution of distance covariance $\mathcal{V}$ Székely et al. (2007). The test statistic is $V\equiv\frac{n\times\mathcal{V}^{2}(X,Y)}{S_{2}}$, where $n$ is the size of smaple and $S_{2}\equiv\frac{1}{n^{2}}\sum_{k,l}|X_{k}-X_{l}|\frac{1}{n^{2}}\sum_{k,l}|Y_{k}-Y_{l}|$. The statistical independence would be rejected if $\sqrt{V}>\Phi^{-1}(1-\alpha/2)$ for selected significance level $\alpha$, where $\Phi$ is the Gaussian cumulative density function.
@@
Figure 5: The statistic $\sqrt{V}$ of signal and background separately for various discriminators. The error bar comes from 1,000 iterations and sample size 10,000 was used. Please note that the scale of y-axis is arbitrary.
It is interesting that the result becomes more independent for all the cases when the jet substructure variables are included for training. This is because jet substructure variables are useful for discrimination of ggH signal from background but relatively less correlated to $m_{\mu\mu}$ compared to $S_{n_{\text{jet}}}$ variables including information about muon and dimuon system. It is obvious that $m_{\mu\mu}$ and the output of discriminators are enormously correlated for (ordinary) DNN. The independence hypothesis is always rejected for DNN case even under $\alpha<0.05$. The situation is improved in the case of BDT but the uncontrolled impact of $m_{\mu\mu}$ still exist. The BDT trained with jet substructure variables barely pass the test up to $\alpha=0.05$. On the other hand, the output of decorrelated DNN and $m_{\mu\mu}$ is independent up to $\alpha=0.215$, which is the upper limit of significance level given in Székely et al. (2007).
@@ III.2 Decorrelated DNN result in detail
After imposing distance correlation loss, the decorrelated DNN loses its accuracy but becomes independent of $m_{\mu\mu}$. Therefore the $m_{\mu\mu}$ can be used without systematic error after categorization. Now, have a detailed look at the categorization result of decorrelated DNN.
Fig.6 shows the number of events for each category and production channel. The number of signal and background is stacked respectively. The vertical dotted lines are the optimal boundaries.
@@
Table.1, 2 shows the number of signal, portion of each production channel, number of background, $S/(S+B)$ and $S/\sqrt{B}$ for optimized categories without and with jet substructure variables. In Fig.6a and Table.1, the portion of VBF among signal is larger for higher purity category. This tendency stems from the discrimination power of two characteristic jets of VBF. On the other hand, the tendency of growing VBF portion is totally reversed for the discriminators trained with jet substructure variables. In Fig.6b, Table.2, the portion of VBF events are getting smaller for higher signal purity category. This reversal of VBF portion represents the improved discrimination power of ggH signals by exploiting jet substructure variables. This property can be used to reduce errors from analysis done separately for each production channel and improve the precision of ggH channel.
@@
IV Discovery significance
The statistical significance is calculated by the profile likelihood ratio fit to the $m_{\mu\mu}$. We used uniform binning of size $50$ MeV. To incorporate systematic errors, nuisance parameters $\theta$ are included with Gaussian suppression. Since the simulation was done for arbitrary luminosity, the number of events are normalized to that of Sirunyan et al. (2021), $137$fb-1. We used non-integers from the normalization directly to Poisson counts to avoid the subtle risk of getting the over- or under-estimated results from the rounding off. The problem of non-integers for Poisson distribution is not a genuine problem since it will be canceled off by calculating likelihood ratio Cowan et al. (2011).
To compare the result, significance is computed for BDT and decorrelated DNN. We do not compute for ordianry DNN case since it is totally distorted. Please note that since we do not extract the number of events by fitting and use the simulation data directly, the significance computed for BDT do not suffer from the problem in extraction though there is distortion indeed. The true significance will be different but cannot predict the tendency since it largely depend on the prescriptions such as the ‘core-pdf’ method used in Sirunyan et al. (2021). Table.3 shows the AUC and significance. Clearly, significance increases about 0.15 by exploiting ggH ISR characteristic. The slightly difference of AUC and significance between BDT and decorrelated dNN can be interpreted as uncontrolled impact of $m_{\mu\mu}$.
Machine BDT decorrelated DNN
Variables $S_{n_{\text{jet}}}$ $J_{n_{\text{jet}}}$ $S_{n_{\text{jet}}}$ $J_{n_{\text{jet}}}$
significance (1.64) (1.78) 1.61 1.76
AUC 0.72 0.77 0.71 0.76
[…]
V Summary
We applied decorrelation using distance correlation to the Higgs to dimuon discovery from ggH channel. The decorrelation between output of discriminator and $m_{\mu\mu}$ allows us to categorize the events without any systematic error from $m_{\mu\mu}$ correlations. The decorrelated DNN achieved the independence simultaneously for both signal and background events up to significance level $\alpha=0.215$. The jet substructure variables are helpful for improving discovery significance of ggH channel and the significance was improved from $1.61$ to $1.76$.
Decorrelation can be helpful in general for both precision and discovery researches as it can preserve the variable of interest-distribution of events after the cut and get rid of systematic errors from distortion in the process of categorization. In addition, all the variables can be used for categorization without any concern of correlation.
VI Acknowledgements
We thank Wonsang Cho for collaboration at an initial stage of the work and for discussion and comments. This work is supported by NRF of Korea grant, No. 0426-20230001.
Appendix A Training variables
The set of training variables for events with each number of jets $n_{\text{jet}}$ will be summarized. We will denote the variables used by the CMS collaboration without jet substructure variables as $S_{n_{\text{jet}}}$ and variables with jet substructure variables as $J_{n_{\text{jet}}}$. Please note that $n_{\text{jet}}=2$ means events with at least two jets. Therefore $S_{2}$ and $J_{2}$ mean training variables for events with multi-jets, not only events with exactly two jets.
A.1 CMS used variables
The training variables were carefully selected to exclude $m_{\mu\mu}$ information. The set of training variables $S_{n_{\text{jet}}}$ used by CMS collaboration are
$~{},\frac{p_{T}^{\mu 1}}{m_{\mu\mu}},\frac{p_{T}^{\mu 2}}{m_{\mu\mu}}\}$
$S_{1}\equiv$ $~{}S_{0}\cup\{p_{T}(j_{1}),\eta(j_{1}),\Delta R(\mu\mu,j_{1})\}$
$S_{2}\equiv$ $~{}S_{1}\cup\{p_{T}(j_{2}),\eta(j_{2}),m_{j_{1}j_{2}},\Delta\eta_{j_{1}j_{2}},\Delta\phi_{j_{1}j_{2}}$
$~{},z_{*},n\}$
In the above notation, CS means those computed in the dimuon Collins-Soper rest frame, $j_{1,2}$ respectively denotes the highest and second highest $p_{T}$ jet, $z*\equiv\frac{y_{\mu\mu}-(y_{j_{1}}+y_{j_{2}})/2}{|y_{j_{1}}-y_{j_{2}}|}$ is the Zeppendfeld variable and $n$ is the number of jets with $p_{T}>25$GeV, $|\eta|<4.7$.
A.2 Jet substructure variables
To exploit the ISR characteristic of ggH channel, we used jet substructure variables in addition to the conventional variables. The jet substructure variables are generally defined as weighted moment sum of constituent transverse momenta normalized by transverse momentum of the jet to capture the fatness of gluon jet. For each jet, we used the number of charged tracks in the jet $n_{\text{track}}$allicchio and Schwartz (2011), girth (linear radial moment) Gallicchio and Schwartz (2011)Gallicchio et al. (2011), broadeningCatani et al. (1992), Energy-energy correlation(EEC) $C^{\beta=0.2}_{1}$ with $\beta=0.2$ Larkoski et al. (2013)Dasgupta et al. (2013), RMS-$p_{T}$allicchio and Schwartz (2011), Pull-vectorGallicchio and Schwartz (2010).
$~{}~{}=\frac{1}{\sum_{i}|\vec{p\ }^{i}|}\displaystyle\sum|\vec{k\ }^{i}_{T}|$
Appendix B Discrimination models in detail
B.1 BDT
w/o jet sub. w/ jet sub.
Variables $S_{0}$ $S_{1}$ $S_{2}$ $J_{1}$ $J_{2}$
$n_{\text{tree}}$ 80 1500 1000 1500 1000
max depth 3 5 7 7 9
learning rate 0.4 0.2 0.2 0.1 0.1
MSL 0.001 0.01 0.02 0.001 0.001
accuracy 0.66 0.65 0.66 0.72 0.73
B.2 (Ordinary) DNN
Table 5 shows the resultant (hyper-)parameters and accuracy for each ordinary DNN machine. Through all the machines, 256 nodes per layer and 100 batches are commonly used.
w/o jet sub. w/ jet sub.
Variables $S_{0}$ $S_{1}$ $S_{2}$ $J_{1}$ $J_{2}$
$n_{\text{layers}}$ 2 3
learning rate 0.007 0.01 0.003 0.04 0.005
epochs 100 300 400 300 300
accuracy 0.76 0.75 0.75 0.79 0.79
$\mathcal{R}$(sig) 0.35 0.39 0.33 0.32 0.24
$\mathcal{R}$(bkg) 0.31 0.31 0.30 0.28 0.18
B.3 Decorrelated DNN
Table 6 shows the resultant (hyper-)parameters and accuracy for each decorrelated DNN machine. Through all the machines, 128 nodes per layers and 100 batches are commonly used. The relative multiplier $\lambda=8\times 10^{-3}$ is also common. We made the $\lambda$ to be magnified to 100 times of its input value if $\mathcal{L}_{\text{decor}}>10^{-3}$ during the training. Therefore the $\lambda=8\times 10^{-1}$ for almost all the time during the training. Since the $\mathcal{L}_{\text{decor}}$ is the sum of distance correlations of signal and background and it is computed with smaller sample, it is larger than the value of distance correlation computed for the test sample(**).
w/o jet sub. w/ jet sub.
Variables $S_{0}$ $S_{1}$ $S_{2}$ $J_{1}$ $J_{2}$
$n_{\text{layers}}$ 3 3 6 3 6
learning rate 0.0004 0.001 0.01 0.001 0.001
epochs 30 40 60 60 60
$\mathcal{L}_{\text{decor}}$* 0.046 0.047 0.044 0.037 0.038
accuracy 0.65 0.64 0.65 0.71 0.72
$\mathcal{R}$(sig)** 0.006 0.009 0.011 0.008 0.008
$\mathcal{R}$(bkg)** 0.019 0.018 0.022 0.010 0.010
Table 6: Resultant model (hyper-)parameters and accuracy. *: $\mathcal{L}_{\text{decor}}$ is the sum of distance correlation loss between score and $m_{\mu\mu}$ respectively for signal and background. 1% of every batch is ampled to calculate $\mathcal{L}_{\text{decor}}$. **: Here the distance correlation $\mathcal{R}$ is calculated with 25,000 events each from the test sample.
Appendix C Independence test
We checked the performance of test using $V$ with the well-known distribution, the two-dimensional Gaussian with various correlation. Fig.7 shows the fraction of rejected samples among 500 iterations(equivalent to type 1 error rate for zero correlation case and statistical power for non-zero correlation cases) with different sample size and correlation. More than sample size 5000, the samples with correlation $\geq$ 0.1 are all rejected for significance level $\alpha\geq$ 0.1.333In Table 1 of Székely et al. (2007), they reported type 1 error rate about 0.1 for various independent distributions, sample sizes under $\alpha=0.1$. In Fig. 2, the power of test for Gaussian with correlation $\rho=0.1$ and under $\alpha=0.1$ is also given.
(a) $\alpha=0.1$
(b) $\alpha=0.215$
(c) $\alpha=0.215$
Figure 7: Fraction of rejected samples among 500 iterations with two-dimensional Gaussian, for various sample size, correlation and significance level $\alpha$
Thus we conducted the test of independence with size 5000 and significance level $\alpha=0.215$ for 500 iterations. The rejected fraction is given in Table.7
Machine BDT DNN decor.DNN
Variable $S_{n_{\text{jet}}}$ $J_{n_{\text{jet}}}$ $S_{n_{\text{jet}}}$ $J_{n_{\text{jet}}}$ $S_{n_{\text{jet}}}$ $J_{n_{\text{jet}}}$
sig(%) 100 76.6 100 100 30.2 17.4
bkg(%) 99.8 64.6 100 100 12.0 16
Table 7: Rejection rate among 500 iterations for the test of dependence. Significance level $\alpha=0.215$ and size = 5000 was used.

### v2

@@ III.1 Comparison
As previously discussed, the decorrelation loss introduced in the decorrelated DNN inherently conflicts with the standard discrimination loss. This trade-off is crucial when comparing the three machine discriminators: BDT, ordinary DNN and decorrelated DNN. We can evaluate them based on two key properties: their discrimination performance (ability to distinguish signal from background) and their degree of decorrelation to the invariant mass $m_{\mu\mu}$.
Fig. 4 displays the Receiver Operating Characteristic (ROC) curve with Area Under the Curve (AUC) for each discriminator configuration. For all three discriminators, including jet substructure variables (solid lines) improves the discrimination performance of ggH channel.
Figure 4: Receiver operating characteristic (ROC) curve for three discrimination schemes - BDT, (ordinary) DNN and decorrelated DNN - with two different set of variables($S_{n_{\text{jet}}}$ and $J_{n_{\text{jet}}})$. The area under curve (AUC) is also given. Note that the true positive rate is equivalent to signal efficiency and the false positive rate is equivalent to (1-background rejection rate).
The ordinary DNN (black line) achieves a substantially higher AUC, which is around 10% higher, compared to BDT (blue line) and decorrelated DNN (red line). This is likely because the ordinary DNN can internally exploit $m_{\mu\mu}$ for discrimination, even though it was not explicitly included as a training variable. However, this high performance comes at a cost of the robustness of analysis related to $m_{\mu\mu}$ (Fig. 2).
In contrast, the BDT and decorrelated DNN, which are inherently limited in their ability to reconstruct $m_{\mu\mu}$ internally, are trained to be insensitive to it. They result in the similar $m_{\mu\mu}$ distributions across score categories (Fig. 1 and 3), but with lower AUCs compared to the ordinary DNN. While both exhibiting similar AUCs, the slight difference in performance between the BDT and decorrelated DNN indicates that there might still be some residual dependence on $m_{\mu\mu}$ in BDT.
The degree of decorrelation between $m_{\mu\mu}$ and the output of discriminators can be seen directly from the distortion of invariant mass distribution categorized by the output of discriminators (Fig. 1, 2, 3). It can also be quantitatively checked by a test of independence using asymptotic distribution of distance covariance $\mathcal{V}$ Székely et al. (2007). The test statistic is
$V\equiv\frac{n\times\mathcal{V}^{2}(X,Y)}{S_{2}}$
where $n$ is sample size and
$S_{2}\equiv\frac{1}{n^{2}}\sum_{k,l}|X_{k}-X_{l}|\frac{1}{n^{2}}\sum_{k,l}|Y_{k}-Y_{l}|$
The statistical independence would be rejected if
$\sqrt{V}>\Phi^{-1}(1-\alpha/2)$
for selected significance level $\alpha$, where $\Phi$ is the Gaussian cumulative density function.
@@
Figure 5: The test statistic $\sqrt{V}$ of signal and background separately for various discriminators. The error bar comes from 1,000 iterations and sample size 10,000 was used. Two lines according to significance level $\alpha=0.05$ and $\alpha=0.215$ are given. Please note that the scale of y-axis is arbitrary.
It is obvious that $m_{\mu\mu}$ and the output of discriminators are enormously correlated for (ordinary) DNN. The independence hypothesis is always rejected for DNN case even under $\alpha<0.05$. The situation is improved in the case of BDT but the uncontrolled impact of $m_{\mu\mu}$ still exist. The BDT trained with jet substructure variables barely pass the test up to $\alpha=0.05$. On the other hand, the output of decorrelated DNN and $m_{\mu\mu}$ is independent up to $\alpha=0.215$, which is the upper limit of significance level given in Székely et al. (2007).
It is interesting that the results become more independent for both BDT and DNN cases when the jet substructure variables are included for training. This is because jet substructure variables have discrimination power other than the $m_{\mu\mu}$ so that the distances of discrimination score have a more spread distribution. The decorrelated DNN result trained with $S_{n_{\text{jet}}}$ is already decorrelated to $m_{\mu\mu}$ and thus there is no further decorrelation by including jet substructure variables.
In summary, it is verified that for all the machine discriminators, the discrimination performance is improved by including jet substructure variables. The gap between AUC of decorrelated DNN and those of other methods is due to the discriminatory power directly attributable to $m_{\mu\mu}$. At the same time, the decorrelation up to $\alpha=0.215$ was achieved by our decorrelation method both choice of training variables, $S_{n_{\text{jet}}}$ and $J_{n_{\text{jet}}}$. It achieves sufficient independence from the invariant mass, for both the Higgs signal and the background simultaneously. The method has a significant potential for more robust Higgs boson to dimuon decay searches and precision measurements using the invariant mass projection.
@@ III.2 Decorrelated DNN result in detail
Imposing distance correlation loss reduces the accuracy of decorrelated DNN but achieves the desired independence from $m_{\mu\mu}$. Consequently, the $m_{\mu\mu}$ can now be reliably used for analysis after the categorization process. We will now delve into a detailed examination of the categorization results of decorrelated DNN.
Fig. 6 shows the number of events for each category and production channel. The number of signal and background is stacked respectively. The vertical dotted lines are the optimal boundaries. Exact number of signal events, background events, $S/\sqrt{B}$ and other information for each optimized category with and without jet substructure variables can be checked in App. D.
@@
In Fig. 6a, the proportion of VBF events within the signal increases for categories with higher purity. On the other hand, the tendency of growing VBF portion is totally reversed for the discriminators trained with jet substructure variables. In Fig. 6b, the portion of VBF events are getting smaller for higher signal purity category. This reversal of VBF portion represents the improved discrimination power of ggH signals by exploiting jet substructure variables. This property allows for the reduction of errors arising from separate analysis of each production channel, ultimately improving the precision of the ggH channel measurement.
IV Discovery significance
We calculated the statistical significance using the profile likelihood ratio fit to the $m_{\mu\mu}$ distribution. We employed uniform binning of size $50$ MeV. To account for systematic errors, nuisance parameters with Gaussian priors were incorporated into the fit. Since the simulation was done for an arbitrary luminosity, the number of events are normalized to that of Sirunyan et al. (2021), $137$fb-1. We used non-integers from the normalization directly to Poisson counts to avoid the subtle risk of getting the over- or under-estimated results from the rounding off. The problem of non-integers for Poisson distribution is not a genuine problem since it will be canceled off by calculating likelihood ratio Cowan et al. (2011).
We computed the statistical significance using both BDT and decorrelated DNN to compare their results. The ordinary DNN was excluded due to significant distortions in the cateogirzed $m_{\mu\mu}$ distribution. Please note that the significance from BDT is also computed under the assumption of statistical independence, which is not true actually. This distortion can give vastly wrong number of events if we fit to the $m_{\mu\mu}$, but our result does not reflect it since we used the simulation data directly rather than extraction from the fitting. The correct computation requires cumbersome combining procedure for significance of correlated categories. We can not predict the significance accurately since it largely depends on the prescriptions such as the ‘core-pdf’ method used in Sirunyan et al. (2021).
Table 1 shows the AUC and significance. Clearly, significance increases about 0.15 by exploiting ggH ISR characteristic. The slight difference of AUC and significance between the BDT and decorrelated DNN can be interpreted as uncontrolled impact of $m_{\mu\mu}$. It should be emphasized that the decorrelated DNN does not necessarily result in increased discovery significance. Decorrelation is an essential ingredient for the analysis with high precision rather than an optional choice.
Machine BDT decorrelated DNN
Variables $S_{n_{\text{jet}}}$ $J_{n_{\text{jet}}}$ $S_{n_{\text{jet}}}$ $J_{n_{\text{jet}}}$
significance (1.64) (1.78) 1.61 1.76
AUC 0.72 0.77 0.71 0.76
[…]
V Summary
We applied decorrelation using distance correlation to the Higgs to dimuon discovery from ggH channel. The decorrelation between output of discriminator and $m_{\mu\mu}$ allows us to categorize the events without any systematic error from $m_{\mu\mu}$ correlations. The decorrelated DNN achieved the independence for both signal and background events simultaneously up to significance level $\alpha=0.215$. The jet substructure variables are helpful for improving discovery significance of ggH channel and the significance was improved from $1.61$ to $1.76$.
Decorrelation can be helpful in general for both precision and discovery as it can preserve the variable of interest-distribution of events after the cut and get rid of systematic errors from distortion in the process of categorization. In addition, all the variables can be used for categorization without any concern of correlation, simplifying the feature selection process. Decorrelation is an essential ingredient for the analysis with high precision. Application of decorreation to the VBF channel and decorrelation using tree-based discriminators can be done in following research.
VI Acknowledgements
We thank Wonsang Cho for collaboration at an initial stage of the work and for discussion and comments. This work is supported by NRF of Korea grant, No. 0426-20230001 and also by the use of computing resources funded by the NRF, Grant No. 2017R1C1B2011048.
Appendix A Training variables
The set of training variables for events with each number of jets $n_{\text{jet}}$ will be summarized. They were carefully selected to exclude the $m_{\mu\mu}$ information.
A.1 Without jet substructure variables
The set of training variables $S_{n_{\text{jet}}}$ without jet substructure variables, which is the same as the one used by CMS collaboration, are
$~{},\frac{p_{T}^{\mu_{1}}}{m_{\mu\mu}},\frac{p_{T}^{\mu_{2}}}{m_{\mu\mu}}\}$
$S_{1}\equiv$ $~{}S_{0}\cup\{p_{T}(j_{1}),\eta(j_{1}),\Delta R(\mu\mu,j_{1})\}$
$S_{2}\equiv$ $~{}S_{1}\cup\{p_{T}(j_{2}),\eta(j_{2}),m_{jj},\Delta\eta_{j_{1},j_{2}},\Delta\phi_{j_{1},j_{2}}$
$~{},z_{*},n\}$
where $p_{T}$ is transverse momentum, $y$ is rapidity, $\theta,\phi$ are zenith and azimuthal angle respectively, $R$ is the absolute distance in $\eta,\phi$ space, $z*\equiv\frac{y_{\mu\mu}-(y_{j_{1}}+y_{j_{2}})/2}{|y_{j_{1}}-y_{j_{2}}|}$ is the Zeppendfeld variable and $n$ is the number of jets with $p_{T}>25$GeV, $|\eta|<4.7$. The index CS means those computed in the dimuon Collins-Soper rest frame, $j_{1,2}$ respectively denotes the highest and second highest $p_{T}$ jet, $\mu_{1},\mu_{2}$ denote two muons, $\mu\mu$ and $jj$ denote dimuon and dijet of $j_{1},j_{2}$ system respectively.
A.2 Jet substructure variables
The jet substructure variables are generally defined as weighted moment sum of constituent transverse momenta normalized by transverse momentum of the jet to capture the fatness of gluon jet. For each jet, we used the number of charged tracks in the jet $n_{\text{track}}$ Gallicchio and Schwartz (2011), girth (linear radial moment) Gallicchio and Schwartz (2011); Gallicchio et al. (2011), broadening Catani et al. (1992), Energy-energy correlation(EEC) $C^{\beta=0.2}_{1}$ with $\beta=0.2$ Larkoski et al. (2013); Dasgupta et al. (2013), RMS-$p_{T}$ Gallicchio and Schwartz (2011), Pull-vector Gallicchio and Schwartz (2010).
$~{}~{}=\frac{1}{\sum_{i}|\vec{p\ }^{i}|}\displaystyle\sum|\vec{k\ }^{i}_{T}|$
Appendix B Distance correlation
The distance covariance ($\operatorname*{\mathcal{V}}$) and distance correlation ($\operatorname*{\mathcal{R}}$) are measures of both linear and non-linear correlation between two random variables/vectors Székely et al. (2007). The distance covariance is defined as an weighted $L_{2}$ norm between the joint characteristic function and the product of marginal characteristic functions ($\phi$).
[…]
where the weight $dw$ is chosen for $\operatorname*{\mathcal{V}}$ to be scale equivariant and therefore for the distance correlation (which is to be defined soon) to be scale invariant. Moreover, to capture the non-linear correlation, $dw$ is chosen to be nonintegrable. It is clear that $X$ and $Y$ are independent if and only if $\mathcal{V}^{2}(X,Y)=0$ by the property of characteristic function. The distance correlation is a distance variance normalized into a value in $[0,1]$.
$\mathcal{R}^{2}(X,Y)\equiv\frac{\mathcal{V}^{2}(X,Y)}{\sqrt{\mathcal{V}^{2}(X,X)\mathcal{V}^{2}(Y,Y)}}$
The distance covariance and distance correlation are also well defined in the sample space. For notational convenience, define doubly centered matrix as $\bar{A}_{kl}\equiv A_{kl}-A_{k\cdot}-A_{\cdot l}+A_{\cdot\cdot}$ where $A_{kl}=|X_{k}-X_{l}|$ and $\cdot$ represent the mean along that index, e.g. for sample size $n$, $A_{k\cdot}\equiv\frac{1}{n}\sum_{l}A_{kl}$. Similarly define $\bar{B}$ for $Y$. Then the empirical distance covariance is defined as
$\mathcal{V}^{2}_{n}(X,Y)$ $\equiv\sum_{k,l}\bar{A}_{kl}\bar{B}_{kl}=S_{1}+S_{2}-2S_{3}$
where
$S_{1}$ $\equiv\frac{1}{n^{2}}\sum_{k,l}A_{kl}B_{kl}$
$S_{2}$ $\equiv\frac{1}{n^{2}}\sum_{k,l}A_{kl}\frac{1}{n^{2}}\sum_{k,l}B_{k,l}=A_{\cdot\cdot}B_{\cdot\cdot}$
$S_{3}$ $\equiv\frac{1}{n^{3}}\sum_{k,l,m}A_{kl}B_{km}$
The empirical distance correlation is defined by the above definition of empirical distance covariance
$\mathcal{R}^{2}_{n}(X,Y)\equiv\frac{\mathcal{V}^{2}_{n}(X,Y)}{\sqrt{\mathcal{V}^{2}_{n}(X,X)\mathcal{V}^{2}_{n}(Y,Y)}}$
The consistency of definitions in population and sample space can be checked in Székely et al. (2007) and the differentiability comes directly from the definition above. Please note that we used empirical distance covariance and correlation without subscript $n$ since there is no ambiguity.
Appendix C Discrimination models in detail
C.1 BDT
w/o jet sub. w/ jet sub.
Variables $S_{0}$ $S_{1}$ $S_{2}$ $J_{1}$ $J_{2}$
$n_{\text{tree}}$ 80 1500 1000 1500 1000
max depth 3 5 7 7 9
learning rate 0.4 0.2 0.2 0.1 0.1
MSL 0.001 0.01 0.02 0.001 0.001
accuracy 0.66 0.65 0.66 0.72 0.73
C.2 (Ordinary) DNN
Table 3 shows the resultant (hyper-)parameters and accuracy for each ordinary DNN machine. 256 nodes per layer and 100 iterations of batch are commonly used.
C.3 Decorrelated DNN
Table 4 shows the resultant (hyper-)parameters and accuracy for each decorrelated DNN machine. 128 nodes per layers and 100 batches are commonly used. In this analysis, we adopted a relative multiplier value of $\lambda=8\times 10^{-3}$. During the training, we dynamically adjust the relative multiplier $\lambda$. If the $\mathcal{L}_{\text{decor}}$ exceeded a threshold of $10^{-3}$, $\lambda$ was magnified by a factor of 100. Since the $\mathcal{L}_{\text{decor}}$ barely touch the threshold, $\lambda$ remained around $8\times 10^{-1}$ for most of the training process. The the sum of distance correlations of signal and background $\mathcal{L}_{\text{decor}}$ is larger than the value of distance correlation computed for the test sample (**) since it is computed with smaller sample.
w/o jet sub. w/ jet sub.
Variables $S_{0}$ $S_{1}$ $S_{2}$ $J_{1}$ $J_{2}$
$n_{\text{layers}}$ 2 3
learning rate 0.007 0.01 0.003 0.04 0.005
epochs 100 300 400 300 300
accuracy 0.76 0.75 0.75 0.79 0.79
$\mathcal{R}$(sig) 0.35 0.39 0.33 0.32 0.24
$\mathcal{R}$(bkg) 0.31 0.31 0.30 0.28 0.18
w/o jet sub. w/ jet sub.
Variables $S_{0}$ $S_{1}$ $S_{2}$ $J_{1}$ $J_{2}$
$n_{\text{layers}}$ 3 3 6 3 6
learning rate 0.0004 0.001 0.01 0.001 0.001
epochs 30 40 60 60 60
$\mathcal{L}_{\text{decor}}$* 0.046 0.047 0.044 0.037 0.038
accuracy 0.65 0.64 0.65 0.71 0.72
$\mathcal{R}$(sig)** 0.006 0.009 0.011 0.008 0.008
$\mathcal{R}$(bkg)** 0.019 0.018 0.022 0.010 0.010
Table 4: Resultant model (hyper-)parameters and accuracy. *: $\mathcal{L}_{\text{decor}}$ is the sum of distance correlation loss between score and $m_{\mu\mu}$ respectively for signal and background. 1% of every batch is ampled to calculate $\mathcal{L}_{\text{decor}}$. **: The distance correlation $\mathcal{R}$s are calculated with 25,000 events each from the test sample.
Appendix D Tables of decorrelated DNN result in detail
Table 5, 6 shows the number of signal, portion of each production channel, number of background, $S/(S+B)$ and $S/\sqrt{B}$ for optimized categories without and with jet substructure variables.